OpenStudy (anonymous):

how do you write 32 ^-3/5 in radical form?? any one know??

OpenStudy (anonymous):

the denominator is the root, the numerator is the power, and the minus sign means take the reciprocal

OpenStudy (anonymous):

the fifth root of 32 is 2, because \(2^5=32\) then you square it and get \(2^2=4\) then you take the reciprocal and get \(\frac{1}{4}\) so all in all \[32^{-\frac{2}{5}}=\frac{1}{32^{\frac{2}{5}}}=\frac{1}{2^2}=\frac{1}{4}\]

OpenStudy (anonymous):

is 1/4 in radical form? I thought radical form meant in likie the sq. rt. sigh?

OpenStudy (anonymous):

sign